OpenStudy (anonymous):

Find an equation of the tangent line to the graph f(x)=2-7x^2 at (2,-26)

OpenStudy (amistre64):

what are 2 things we need to know to define a line?

OpenStudy (anonymous):

two points?

OpenStudy (amistre64):

well, 2 points are useful yes, but we only have one point given to us ... what else could we use to define a line with? a point and ....

OpenStudy (anonymous):

slope?

OpenStudy (amistre64):

correct, so a point and a slope can be used to define a line. its called the point slope form of the equation. now, how do we determine the slope of a curve at a given point?

OpenStudy (anonymous):

the change in y over the change in x? so the point slope equation?

OpenStudy (amistre64):

hmm, im assuming you know what a derivative is for this to work.

OpenStudy (anonymous):

yeah but I don't really get derivatives

OpenStudy (anonymous):

I know there's 4 types

OpenStudy (amistre64):

you should have a simple rule you can apply constant rule, and a power rule

OpenStudy (amistre64):

\[\frac{d}{dx}k=0\] \[\frac{d}{dx}x^n=n~x^{n-1}\]

OpenStudy (anonymous):

I'm just having a hard time applying the equation to these rules

OpenStudy (amistre64):

9 - x^2 ^ ^ ^ this is a power of x ^ and this is a constant

OpenStudy (amistre64):

what is the derivative of 9? what is the derivative of -x^2?

OpenStudy (amistre64):

e3rr, -7x^2 ... forgot the equation we were dealing with :/

OpenStudy (anonymous):

so for 2 it would be 0 and -7x^2 would be -7x^-6?

OpenStudy (amistre64):

\[\frac{d}{dx}(2-7x^2)\] \[\frac{d}{dx}(29)-\frac{d}{dx}(7x^2)\] \[\frac{d}{dx}(2)-7\frac{d}{dx}(x^2)\] \[0-7(2x^{2-1})\]

OpenStudy (anonymous):

how did you get 29?

OpenStudy (amistre64):

typo, slip of the fingers.

OpenStudy (anonymous):

can you please explain how you got the answer step by step?

OpenStudy (amistre64):

i jsut did ... except for 29 which is spose to be a 2 instead use the constant rule for the constant, and the power rule for the power of x there is also a 'pull out' rule which lets us pull out that 7 to the outside.

OpenStudy (amistre64):

your -7x^2 derivative is off ... try it again

OpenStudy (amistre64):

the format for the tangent line is just gonna be: y - f(a) = f'(a) (x-a) so as long as we can determine f'(a) we are set

OpenStudy (anonymous):

sorry my book shows different explanations so I'm a little confused right now

OpenStudy (amistre64):

tell me your thoughts then, what does your book explain?

OpenStudy (anonymous):

it basically shows me all these rules but all these variables are confusing me

OpenStudy (anonymous):

but the examples only show me if I have a point and slope

OpenStudy (amistre64):

what is your rule for powers of x?

OpenStudy (amistre64):

can you take a picture of it?

OpenStudy (anonymous):

For any real number k if y=x^k then d/dx(x^h)= k*x^k-1

OpenStudy (amistre64):

let k=2 x^2 become 2x^(2-1) or simply 2x

OpenStudy (amistre64):

so our derivative is -7x^2 derives to -7(2x), at x=2 that is ... -7(4)

OpenStudy (anonymous):

so the 2 becomes x?

OpenStudy (amistre64):

out point (2,-26) defines the value of x as 2

OpenStudy (amistre64):

*our point ...

OpenStudy (amistre64):

2x, when x=2 is just ... 2(2) right?

OpenStudy (anonymous):

how did x^2 become 2x?

OpenStudy (amistre64):

by the derivative rule you posted. let k=2

OpenStudy (amistre64):

x^k derives to k*x^(k-1)

OpenStudy (anonymous):

Ohhhh I get it now

OpenStudy (amistre64):

by first principles of limit \[\lim_{h\to0}\frac{(x+h)^2-x^2}{(x+h)-x}\] \[\lim_{h\to0}\frac{x^2+h^2+2xh-x^2}{h}\] \[\lim_{h\to0}\frac{h^2+2xh}{h}\] \[\lim_{h\to0}\frac{h(h+2x)}{h}\] \[\lim_{h\to0}(h+2x)=2x\]

OpenStudy (anonymous):

but I'm still pretty confused

OpenStudy (amistre64):

the 'rules' are really a table of outcomes that can be applied so that you dont have to use the long limiting process each and every time.

OpenStudy (anonymous):

so now that we know the derivative of -7x^2 to be -7(2x) which is -7(2*2)= -28 what do we do next?

OpenStudy (amistre64):

well, thats our slope of our tangent line. the rest is from algebra class.

OpenStudy (amistre64):

construct the line using your slope and the stated point

OpenStudy (anonymous):

y=mx+b

OpenStudy (anonymous):

or point slope equation y-y1=m(x-x1)

OpenStudy (amistre64):

y - f(a) = f'(a) (x-a) or rewritten as y = f'(a) (x-a) +f(a)

OpenStudy (amistre64):

point slope equation

OpenStudy (anonymous):

y-(-26)=-28(x-2)

OpenStudy (amistre64):

thatll do it

OpenStudy (anonymous):

y=-28x+30

OpenStudy (anonymous):

could I get a two out of it to be 2(-14x+15)? or is okay like that?

OpenStudy (amistre64):

usually we dont factor it, keep it in y=mx+b form is conventional

OpenStudy (anonymous):

so that's my final answer? thank you! is there any way I reward you on here? it's my first time using this site

OpenStudy (amistre64):

there should be a 'best response' button to everything ive posted. clicking it gives out a medal. but saying thank you and learning a little bit is reward enough for me to be honest.

OpenStudy (anonymous):

okay thank you!!

OpenStudy (amistre64):

youre welcome :) and good luck